2/3a +5 = 1/3 (5a +3)
Answer by checkley77(12844)

2/3a +5 = 1/3 (5a +3)
2a/3+5=(5a+3)/3
2a/3-(5a+3)/3=-5
(2a-5a-3)/3=-5
(-3a-3)/3=-5
-a-1=-5
-a=-5+1
-a=-4
a=4 ans.
Proof:
2*4/3+5=(5*4+3)/3
8/3+5=(20+3)/3
8/3+15/3=23/3
23/3=23/3